AIR-BROADENED LINE WIDTHS AND SHIFTS IN THE ν_{3} BAND OF ^{16}O_{3} AT TEMPERATURES BE-TWEEN 160 AND 300 K

\underline{M. A. H. SMITH}, C. P. RINSLAND, Science Directorate, NASA Langley Research Center, Hampton, VA 23681-2199; V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795; and A. M. COX, New Horizons Governor’s School for Science and Technology, Hampton, VA 23666 and Poquoson High School, Poquoson, VA 23662.

The 9.6-μm bands of O_{3} are used by many remote-sensing experiments for retrievals of terrestrial atmospheric ozone concentration profiles. Line parameter errors can contribute significantly to the total errors in these retrievals, particularly for nadir-viewing^{a}. We have used the McMath-Pierce Fourier transform spectrometer at the National Solar Observatory on Kitt Peak to record numerous high-resolution infrared absorption spectra of O_{3} broadened by various gases at temperatures between 160 and 300 K. Over 25 air-broadened spectra were analyzed simultaneously using a multispectrum nonlinear least squares technique^{b} to determine Lorentz pressure-broadening and pressure-induced shift coefficients along with their temperature dependences for selected P- and R-branch transitions in the ν_{3} fundamental band of ^{16}O_{3}. We have compared the present results with other measurements reported in the literature for O_{3} broadened by air or by N_{2}.

aJ. Worden, S. S. Kulawik, M. W. Shephard, S. A. Clough, H. Worden, K. Bowman and A. Goldman, J. Geophys. Res. \underline{109}, 9308-9319 (2004).

bD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).